OpenStudy (anonymous):

A HYPOTENUSE OF A 45 45 90 TRIANGLE HAS A LENGTH OF 10 UNITS. WHAT IS THE LENGTH OF ONE OF ITS LEGSHELP!!

OpenStudy (solomonzelman):



OpenStudy (solomonzelman):

Don't forget to rationalize the denominator.

OpenStudy (anonymous):

...

OpenStudy (anonymous):

that makes no sense

OpenStudy (anonymous):

do you know the pythagorean theorem?

OpenStudy (anonymous):

yes a^2+b^2=c^2

OpenStudy (anonymous):

i know c^2 =10

OpenStudy (solomonzelman):

These are the proportions for 45-45-90 triangle (proven by the Pythagorean theorem) You just need to divide the hypotenuse by `√2`

OpenStudy (anonymous):

also because the triangle is a 45 45 90 triangle the legs are equal. therefore, a=b in the pythagorean thereom.

OpenStudy (anonymous):

7.07?

OpenStudy (anonymous):

would it be 10 radical 2

OpenStudy (anonymous):

7.07 or 5 radical 2

OpenStudy (solomonzelman):

No .... !!! Never approximate √2, your teachers will lower your grade SO MUCH because of that. Guarantee!

OpenStudy (anonymous):

@nabiladh you look hot btw ;)

OpenStudy (anonymous):

the only choices are a) 10 rad 2 b)5 rad 2 c)5 unit d)10 unit

OpenStudy (anonymous):

@BestRivenNA hahahaha r u 12

OpenStudy (anonymous):

its b)

OpenStudy (anonymous):

12 as in 12 years old?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ummm no?